Parth (parthkohli):

``` Prove using the Division Algorithm that an integer is either even or odd, but never both. ``` Is mine a legitimate proof?

Parth (parthkohli):

Okay, so I did this: A number is either \(0\pmod{2}\) or \(1\pmod{2}\). Using the division algorithm, we may write both statements as \(n = 2k + 0\) or \(n = 2k + 1\) and \(k\in\mathbb{Z}\). As seen, these two are just the definitions of even and odd numbers.

Parth (parthkohli):

@AccessDenied Is there a theorem that tells us that a number is either \(0\pmod{2}\) or \(1 \pmod{2}\)?

Parth (parthkohli):

Or roughly, in the statement \(x\pmod{y}, \ \ x,y\in\mathbb{N}\), there are \(y\) number of possibilities of \(x\)?

OpenStudy (anonymous):

What is the idea behind using the scroll bar there above..??

Parth (parthkohli):

It made a scroll bar automatically.

OpenStudy (anonymous):

Really?? Without using any code or command??

Parth (parthkohli):

I intended to put it inside a quotation block.

Parth (parthkohli):

Anyway, how do I go about proving that a number can never be both?

OpenStudy (anonymous):

NO GOOD AT IT..

OpenStudy (anonymous):

``` Parth Kohli ```

OpenStudy (anonymous):

Just trying how you did that..